# On the excitability of two-level atoms by spectrally encoded single-photon wave packets

### v1

@@ Abstract
Single-photon wave packets are utilized as quantum information carriers within quantum information science and quantum communication applications. Here, we analyze the time-dependent interaction of a two-level atom with a spectrally encoded single-photon wave packet and obtain the excitation probability of the atom based on the Heisenberg-Langevin equations. Our results show that the spectral encoding process causes the single-photon wave packet to spread in the time domain, and the intensity of the wave packet decreases. Therefore, an encoded single-photon wave packet excites a two-level atom much weaker than an uncoded one. These results are crucial for understanding how encoded quantum light interacts with quantum nodes in realistic quantum networks. In particular, they reveal limitations and trade-offs in using spectrally encoded photons for secure, multiplexed quantum communication. The findings offer insights into optimizing encoding strategies for efficient atom-photon interactions and contribute to the development of scalable and secure quantum networking protocols.
keywords: Atom-photon interaction, Heisenberg-Langevin equations, Spectral encoding, Photon wave packet.
@@ I Introduction
The paper is organized as follows. In section II, we review a general formalism to analyze the dynamics of a two-level atom that interacts with a continuous mode state of quantum light. In section III, first, we introduce the spectral encoding approach. Then, we investigate the dynamics of interaction between a two-level atom with an encoded single-photon wave packet to find the excitation probability of the atom. Finally, the results are discussed in section IV.
@@ II.1 Excitation of a two-level atom by a rectangular single-photon wave packet
$\xi(\omega)=\begin{cases}\sqrt{P_{0}}/W&\mbox{for}-\frac{W}{2}\leq\omega\leq\frac{W}{2}\\ 0&\mbox{elsewhere}\end{cases}$ (9)
@@
where sinc x=sin $\mbox{sinc x}=\frac{\mbox{sin x}}{\mbox{x}}$.
@@ III.2 Excitation of a two-level atom by an encoded single-photon wave packet
$\xi^{e}(t)=\frac{\sqrt{P_{0}}}{N_{0}}\mbox{sinc}\big{(}\frac{\Omega}{2}t\big{)}\sum_{n=-N}^{N}\mbox{exp}\big{[}-i(n\Omega t)+\phi_{n}\big{]}$ (15)
where the phase $\phi_{n}$ denotes the $n$th element of a code consisting of $N_{0}=2N+1$. Moreover, the term $\mbox{sinc}\big{(}\frac{\Omega}{2}t\big{)}$ is a real function that represents the temporal width of the encoded wave packet and the exponential term shows a periodic behavior with period $T=2\pi/\Omega$ which its temporal shape depends on the code elements $\phi_{n}$. Note that if we have $\phi_{n}=0$ for all $n$, Eq. (15) reduces to Eq. (10) Salehi .
We now consider $N_{0}=3,5,7,31,63$ for the code length of an encoded single-photon wave packet to calculate the excitation probability $P_{e}(t)$ of a two-level atom which interacts with an encoded pulse according to Eq. (7). In addition, we assume the value of the phase $\phi_{n}$ is 00 and $\pi$ which correspond to the amplitude of $+1$ and $-1$, respectively. This spectral encoding process is called random binary spectral encoding Salehi .
@@
Figure 5: The excitation probability $P_{e}(t)$ of a two-level atom by an encoded photon wave packet considering three different sequences of the random phase (0 and $\pi$) for the code length $N_{0}=63$ in the binary spectral encoding process.
@@
Figure 7: Spreading the encoded photon wave packet in time domain for the various code lengths. a) $N_{0}=3$ b) $N_{0}=5$ c) $N_{0}=7$ d) $N_{0}$=31 and e) $N_{0}=63$.
IV Conclusion
We studied the time-dependent interaction of a two-level atom with a spectrally encoded single-photon wave packet based on the Heisenberg-Langevin equations. We used a spectral phase encoding method inspired by classical CDMA systems to encode quantum information in the frequency domain of the photon. Considering this formalism, we showed that for an encoded wave packet, the excitation probability of the atom is much lower than an unencoded one. On the other hand, the encoding process causes the wave packet to spread in the time domain and the ability to excite a two-level atom decreases. Moreover, we evaluated how encoding parameters—such as code length and phase randomness—affect atomic excitation probabilities.
These findings are particularly relevant for quantum networks, where atoms serve as quantum nodes and qubits are transmitted over long distances via single-photon wave packets. In such systems, an encoded single-photon wave packet carries quantum information to the receiver. If the receiver applies a decoding mask that is the conjugate of the original encoding mask, the photon’s spectral profile is correctly reconstructed, allowing it to interact effectively with the atomic node. However, if the decoding mask does not match the encoding mask, the wave packet fails to recover its original shape and remains a temporally spread, low-intensity pulse. As a result, it has minimal impact on the atomic excitation. This behavior provides a built-in filtering mechanism: atoms can act as selective photodetectors, responding only to correctly encoded photons while rejecting mismatched or unauthorized signals. Such properties are valuable for enhancing the security and robustness of quantum communication protocols.
Moreover, this research provides a fundamental insight into how spectral encoding, commonly used in CDMA systems, affects the quantum mechanical behavior of light-matter interaction. Understanding this effect is essential for developing robust quantum communication protocols that rely on encoding information in the frequency domain. Our findings indicate that while spectral encoding enhances information security and multiplexing capabilities, it also introduces trade-offs in terms of interaction efficiency with quantum nodes, which must be addressed in the system design. This study also opens the door to future research in optimizing spectral encoding schemes that balance encoding complexity and atom-photon interaction strength. In practical implementations of quantum repeaters or quantum routers, such trade-offs must be well-understood to ensure high fidelity and reliability of quantum information transfer. Furthermore, the methodology presented here can be extended to investigate other encoding formats, multi-level atoms, and more complex photonic structures, including entangled wave packets or time-bin encodings.
Ultimately, our work contributes to the theoretical foundation necessary for advancing quantum network architectures and can inform experimental setups aiming to test photon-mediated interactions at the quantum level with encoded information. The ability to predict and control atomic excitation probabilities in the presence of encoded quantum signals is a step toward realizing scalable and secure quantum communication infrastructures.

### v2

@@ Abstract
We analyze the time-dependent interaction between a two-level atom and a spectrally encoded single-photon wave packet using the Heisenberg-Langevin equations and derive the atomic excitation probability. Spectral phase encoding broadens the photon wave packet in the time domain and reduces its peak intensity, leading to markedly weaker atomic excitation than for an unencoded photon. We formalize this behavior via an overlap bound with the time-reversed spontaneous-emission mode and show how excitation depends on code length, bandwidth, and phase errors. Interpreted at the quantum network level, atoms behave as phase-sensitive, and mode-selective receivers whose response scales with a spectral-overlap functional that captures decoding fidelity, detuning, and multiuser interference. From this, we extract design rules and performance bounds for encoded links, quantifying trade-offs among code length, addressability, cross-talk, and identifying tolerances for decoding error. These results clarify how spectrally encoded photons couple to quantum nodes and provide guidelines for efficient, scalable, and secure quantum networking.
keywords: Atom-photon interaction, Heisenberg-Langevin equations, Spectral encoding, Quantum networks.
@@ I Introduction
Moreover, as we know, a central challenge for quantum networks is the selective addressability of remote atomic or solid-state nodes while sharing photonic resources among many users. In this context, spectral encoding provides a high-dimensional multiplexing scheme compatible with optical fibers and ultrafast pulse shaping. However, atom-photon interfaces are inherently mode-selective, as the coupling efficiency depends on the spectral-temporal overlap between the photon wave packet and the atomic or cavity response. In this work, we quantify this overlap for encoded single-photon wave packets and, crucially, translate it into network-level figures of merit-decoding fidelity, multiuser cross-talk, and mismatch tolerance, thereby turning our microscopic model into concrete design guidance for scalable, multiplexed quantum links.
The paper is organized as follows. In section II, we review a general formalism to analyze the dynamics of a two-level atom that interacts with a continuous mode state of quantum light. In section III, first, we introduce the spectral encoding approach. Then, we investigate the dynamics of interaction between a two-level atom with an encoded single-photon wave packet to find the excitation probability of the atom. In section IV, we analyze the applications of our approach in quantum networks. Finally, the results are discussed in section V.
@@ II.1 Excitation of a two-level atom by a rectangular single-photon wave packet
[…]
@@
[…]
@@
II.2 Optimal temporal mode and an overlap bound
It is well known that, for a two-level atom coupled to a propagating mode under the Markov approximation, the time-reversed spontaneous-emission mode maximizes single-photon absorption Stobinska . Denoting the optimal (normalized) temporal profile by
$\xi_{\operatorname{opt}}(t)=\sqrt{\gamma}e^{\gamma t/2}\Theta(-t)$ (11)
where $\Theta(-t)$ is a Heaviside step function, the excitation probability achieved by an arbitrary normalized photon temporal mode $\xi(t)$ obeys the overlap bound
$P_{e}^{\max}\leq\beta|\int_{-\infty}^{\infty}\operatorname{d}t\xi^{*}(t)\xi_{\operatorname{opt}}(t)|^{2}=\beta|\langle\xi\mid\xi_{\operatorname{opt}}\rangle|^{2}$ (12)
where $\beta\in[0,1]$ denotes the coupling efficiency into the atomic radiation mode for free-space focusing or waveguide coupling. In other words, $\beta$ measures the fraction of the atom’s emission that goes into the mode of interest-equivalently, the mode-matching efficiency or branching ratio into that channel. Moreover, we have $\int\operatorname{d}t\mid\xi(t)\mid^{2}=1$. The bound in Eq. (12) follows directly from the Heisenberg-Langevin Eqs. (3)–(4) by writing the atomic drive as a convolution with the impulse response $h(t)=\sqrt{\gamma}e^{-\gamma t/2}\Theta(t)$ and using Cauchy–Schwarz, the equality is saturated for $\xi=\xi_{\operatorname{opt}}$. In our simulations, $\beta=1$ (ideal coupling), so Eq. (12) provides a tight, parameter-free reference for all cases studied. The derivation details of Eq. (12) are given in Appendix A.
@@ III.2 Excitation of a two-level atom by an encoded single-photon wave packet
$\xi^{e}(t)=\frac{\sqrt{P_{0}}}{N_{0}}\mbox{sinc}\big(\frac{\Omega}{2}t\big)\sum_{n=-N}^{N}\mbox{exp}\big[-i(n\Omega t+\phi_{n})\big]$ (17)
where the phase $\phi_{n}$ denotes the $n$th element of a code consisting of $N_{0}=2N+1$. Moreover, the term $\mbox{sinc}\big(\frac{\Omega}{2}t\big)$ is a real function that represents the temporal width of the encoded wave packet and the exponential term shows a periodic behavior with period $T=2\pi/\Omega$ whose temporal shape depends on the code elements $\phi_{n}$. Note that if we have $\phi_{n}=0$ for all $n$, Eq. (17) reduces to Eq. (10).
We now consider $N_{0}=3,5,7,31,63$ for the code length of an encoded single-photon wave packet to calculate the excitation probability $P_{e}(t)$ of a two-level atom which interacts with an encoded pulse according to Eq. (7). In addition, we assume the value of the phase $\phi_{n}$ is 0 and $\pi$ which correspond to the amplitude of +1+1 and 1-1, respectively. This spectral encoding process is called random binary spectral encoding Salehi .
In the binary spectral encoding, we intentionally choose odd $N_{0}$ so that a central chip is centered at $\omega_{0}$ (the atomic transition), yielding a symmetric index set $n\in\{-N,...,0,...,N\}$. This choice avoids placing $\omega_{0}$ in between two chips (the even-$N_{0}$ case), which tends to introduce a systematic $\cos(\Omega t/2)$ modulation in the time-domain envelope and can lead to strong destructive interference at $t=0$ for many random $\{0,\pi\}$. We verified numerically that even $N_{0}$ exhibits a larger fraction of near-vanishing overlaps with $\xi_{\operatorname{opt}}(t)$, consistent with the analytical discussion below. In other words, for even $N_{0}$, the resonance lies between the two central chips $n\pm\frac{1}{2}$. Writing the sum over half-integer indices introduces an extra factor $\cos(\Omega t/2)$ in $c_{\phi}(t)$, which suppresses the integrand near $t\approx 0$ for random $\{0,\pi\}$ phases and thus yields a larger set of near-zero overlaps in Eq. (12).
@@
Figure 5: The excitation probability $P_{e}(t)$ of a two-level atom by an encoded photon wave packet considering three different sequences of the random phase (0 and $\pi$) for the code length $N_{0}=63$ in the binary spectral encoding.
@@
Figure 7: Spreading the encoded photon wave packet in the time domain for the various code lengths. a) $N_{0}=3$ b) $N_{0}=5$ c) $N_{0}=7$ d) $N_{0}$=31 and e) $N_{0}=63$.
Using Eq. (17), the normalized encoded temporal profile can be expressed as a product of a slowly varying envelope $s(t)=\frac{\sqrt{P_{0}}}{N_{0}}\mbox{sinc}(\frac{\Omega t}{2})$ and a coded comb
$c_{\phi}(t)=\sum_{n=-N}^{N}\exp[-in\Omega t+i\phi_{n}]$ (18)
so that $\xi_{e}(t)\propto s(t)c_{\phi}(t)$. The overlap in Eq. (12) becomes
$\langle\xi_{e}\mid\xi_{\operatorname{opt}}\rangle\propto\int\operatorname{d}t\mbox{sinc}(\frac{\Omega t}{2})e^{\gamma t/2}\Theta(-t)\sum_{n=-N}^{N}\exp[in\Omega t-i\phi_{n}]$ (19)
Evaluating the time integral term-wise yields a closed-form dependence on $\Omega/\gamma$ and on the code phases $\{\phi_{n}\}$ via a weighted discrete Fourier transform of the sequence $\{e^{-i\phi_{n}}\}$. Consequently, (i) increasing the code length $N_{0}$ broadens $c_{\phi}(t)$ and reduces the overlap, explaining the monotonic trend of FIG. 6; and (ii) different phase sequences produce distinct beating patterns in $P_{e}$ harmonics of $\Omega$ consistent with the oscillations seen in FIG. 5.
IV Applications to quantum networks: mode-selective atom-photon interfaces with spectral encoding
IV.1 Decoding as spectral-phase compensation
Let the encoder impose a phase mask $\theta(\omega)$, producing the encoded spectrum $\xi^{e}(\omega)=\xi(\omega)e^{-i\theta(\omega)}$. A receiver implements a decoding mask $D(\omega)=e^{i\tilde{\theta}(\omega)}$ (e.g., via a $4f$ pulse shaper), yielding an effective input $\xi_{in}(\omega)=\xi(\omega)e^{i[\tilde{\theta}(\omega)-\theta(\omega)]}$. We define the residual phase error $\phi(\omega)=\tilde{\theta}(\omega)-\theta(\omega)$. So, perfect decoding corresponds to $\phi(\omega)=0$.
The atomic response in the frequency domain is well captured by the Lorentzian susceptibility (narrowband approximation in Section II) Scully
$\chi_{A}(\omega)=\frac{\sqrt{\gamma}}{\gamma/2-i(\omega-\omega_{A})}$ (20)
where $\gamma$ is the dipole decay rate and $\omega_{A}$ is the atomic transition frequency. In the single-photon, weak-excitation regime used throughout, the excitation amplitude is proportional to the scalar product of the input mode with $\chi_{A}$
$\mathcal{A}\propto\int\operatorname{d}\omega\xi(\omega)e^{i\phi(\omega)}\chi_{A}(\omega)$ (21)
We define the spectral-overlap functional Fiechtner
$\mathcal{M}[\phi,\Delta]=\frac{\int\operatorname{d}\omega\xi(\omega)e^{i\phi(\omega)}\chi_{A}(\omega-\Delta)}{\Big(\int\operatorname{d}\omega\mid\xi(\omega)\mid^{2}\Big)^{1/2}\Big(\int\operatorname{d}\omega\mid\chi_{A}(\omega)\mid^{2}\Big)^{1/2}}$ (22)
with dtuning $\Delta=\omega-\omega_{A}$. The spectral-overlap functional was defined by inserting the photon spectrum $\xi(\omega)$ and the atomic susceptibility $\chi_{A}(\omega)$ into the excitation amplitude integral and normalizing it to be dimensionless. It is simply the normalized inner product between the photon mode and the atomic response. This functional directly measures how well the photon’s spectral-temporal mode matches the atomic absorption mode, and the excitation probability satisfies $P_{e}^{\max}\propto\mathcal{M}[\phi,\Delta]\mid^{2}$. Then, the peak excitation probability obeys the bound in Eq. (12)
$P_{e}^{\max}\leq\beta\kappa\mid\mathcal{M}[\phi,\Delta]\mid^{2}$ (23)
where $\beta\in[0,1]$ is the coupling efficiency into the atomic mode (free-space or waveguide) and $\kappa$ is a geometry-dependent constant (absorbed in our normalization in Section II). Equality is approached when temporal mode matching is optimal. This recovers our observed suppression for encoded, undecoded photons, and quantifies recovery under correct decoding $(\phi=0)$.
Bandwidth matching
For our rectangular spectrum of width $W$ in Eq. (9), the perfectly decoded case $(\phi=0)$ yields
$\mid\mathcal{M}[0,\Delta]\mid^{2}\approx\frac{1}{N}\int_{-W/2}^{W/2}\frac{\gamma\operatorname{d}\omega}{\gamma^{2}/4+(\omega-\Delta)^{2}}$ (24)
with $N=\int\operatorname{d}\omega\mid\xi(\omega)\mid^{2}$. This increases with $W/\gamma$ up to the point where the Lorentzian tails limit further gains, for a given $\gamma$ there is an optimal $W$ balancing spectral weight near resonance against total pulse energy. This explains why increasing code length $N_{0}$ (fixed $W$) does not improve excitation unless decoding removes the imposed phase structure.
Our analysis uses a two-level, Markovian model with a single decay rate $\gamma$ and neglects recoil and multi-level structure. In free space, the achievable excitation also depends on the mode-matching parameter $\beta\leq 1$, which multiplies the overlap in Eq. (12). Non-Markovian reservoirs, structured continua, or multi-level $\Lambda$-systems can modify the optimal mode and the decoding strategy. Nonetheless, the overlap-based interpretation carries over: unitary spectral mode conversion preserves inner products, so matched decoding remains the key to recovering strong atom-photon coupling with spectrally encoded single photons.
It is important to note that both encoding and decoding are unitary linear transformations in the single-photon Hilbert space, so they do not constitute a measurement and do not extract information about the encoded state. The process does not involve copying or amplifying a quantum state, rather, it merely rotates the state within its Hilbert space. Therefore, the no-cloning theorem, which forbids the creation of identical, independent copies of an arbitrary unknown state-remains fully respected.
IV.2 Mismatch tolerance and phase-noise budget
We assume small residual phase errors with zero mean and variance $\sigma_{\phi}^{2}$ over the occupied band. A second-order cumulant expansion gives
$\mathcal{M}[\phi,\Delta]\approx\mathcal{M}[0,\Delta]e^{-\sigma_{\phi}^{2}/2}\implies P_{e}\approx P_{e}^{(0)}e^{-\sigma_{\phi}^{2}}$ (25)
So a $10\%$ loss in $P_{e}$ corresponds to $\sigma_{\phi}\approx\sqrt{0.105}\approx 0.32$ rad rms across the in-band frequencies. This sets a concrete specification for pixel-to-pixel SLM (spatial light modulator) errors and spectral-phase drift in a fielded link.
For a binary code with chips of width $\Omega=W/N_{0}$ and random sign errors of probability $p$, one obtains (Appendix B)
$\langle e^{i\phi}\rangle=(1-p)(+1)+p(-1)=1-2p\implies P_{e}\approx P_{e}^{(0)}(1-2p)^{2}$ (26)
so $p\leq 5\%$ keeps $P_{e}\geq 0.81P_{e}^{(0)}$.
IV.3 Multiuser CDMA and cross-talk at an atomic node
Consider $K$ simultaneous users with mutually assigned binary phase codes $\{\theta_{k}(\omega)\}$. The node intends to receive user $j$ and applies a decoding mask $D_{j}(\omega)=e^{i\theta_{j}(\omega)}$. The effective field is a sum $\sum_{k}\xi(\omega)e^{i[\theta_{j}(\omega)-\theta_{k}(\omega)]}$. The desired term has $\phi=0$; the $k\neq j$ terms contribute with code-difference masks. For random $\pm 1$ chip phases with length $N_{0}$, the expected value over random code assignments can be written as
$\mathbb{E}\Big[\lvert\int\operatorname{d}\omega\xi(\omega)e^{i[\theta_{j}-\theta_{k}]}\chi_{A}(\omega)\rvert^{2}\Big]\propto\frac{1}{N_{0}}$ (27)
i.e., cross-talk scales as $N_{0}^{-1}$. Hence, the single-node Signal-to-Interference Ratio satisfies $\text{SIR}\sim(K-1)^{-1}N_{0}$ (up to constants from spectral shaping and $\chi_{A}$). For practical guidance, we can keep $K\lesssim cN_{0}$ with $c<1$ (e.g., $c\sim 0.2-0.3)$ to maintain $\text{SIR}\gtrsim 3-5$ at the atomic interface. This formalizes the multiplexing benefit and the trade-off with code length observed in Section III. For more mathematical details, please refer to Appendix C.
IV.4 Security and addressability via mode selectivity
In the context of quantum communication, mode selectivity provides an inherent mechanism for both secure access and user addressability. The atomic system functions as a natural mode filter whose excitation probability $P_{e}$ depends quadratically on the spectral-temporal overlap between the incoming photon and the atomic response. When a photon arrives with a mismatched spectral-phase pattern, the destructive interference between frequency components suppresses the field amplitude at the atomic resonance. As demonstrated in Section III, the excitation probability then decreases rapidly with increasing code length $N_{0}$, yielding suppression factors that can reach several orders of magnitude for long encoded sequences. From a network standpoint, this behavior acts as a built-in physical layer of security: only photons whose spectral mode precisely matches the decoding mask $D(\omega)$ and the atomic resonance are capable of efficient excitation, while any unauthorized or misaligned signals are effectively rejected by the atom itself. This selectivity simultaneously enables addressability-each atomic node can be assigned a distinct decoding mask, allowing multiple users to share the same optical channel without classical routing. Consequently, the atom-photon interface serves as a passive, noise-free discriminator of spectral codes, enhancing both the privacy and scalability of quantum networks.
IV.5 Co-design guidelines
The quantitative analysis above suggests several practical guidelines for the co-design of spectral encoders, decoders, and atom-photon interfaces.
First, bandwidth matching between the encoded photon spectrum and the atomic linewidth is essential. The excitation probability $P_{e}$ increases with the ratio $W/\gamma$ up to an optimum value where spectral weight begins to fall outside the Lorentzian response of the atom. We proved that for most single-transition systems, maximum excitation occurs for $W/\gamma\approx 1.39$ (Appendix D). Second, the code length $N_{0}$ should be chosen to balance multiplexing capacity and robustness. Longer codes improve the signal-to-interference ratio approximately as $N_{0}$ but they also increase sensitivity to phase errors. The tolerance analysis in Section IV B, shows that maintaining an rms phase error $\sigma_{\phi}\lesssim 0.3$ rad limits excitation loss to below $10\%$, implying a strict calibration requirement on the pulse shaper or spatial light modulator (SLM). Third, maximizing the coupling efficiency $\beta$ between the optical field and the atomic dipole remains the single most effective way to improve performance, as $P_{e}^{\max}\propto\beta\mid\mathcal{M}\mid^{2}$. High-NA (Numerical Aperture) free-space coupling, fiber-cavity, or nanophotonic waveguide geometries can substantially enhance $\beta$. Finally, spectral detuning $\Delta$ should be minimized or pre-compensated in the decoding phase $D(\omega)$ by introducing an appropriate linear phase term to center the photon spectrum on the atomic transition. Together, these guidelines translate the microscopic excitation model into an experimentally actionable design toolkit for encoded quantum-network links.
IV.6 Experimental pathway
The theoretical framework developed in this work can be validated on a variety of experimental platforms using existing photonic and atomic technologies. A representative configuration is as follows.
A single-photon source, for example, spontaneous parametric down-conversion or quantum-dot emission spectrally shaped by a $4f$ pulse-shaping setup equipped with a programmable SLM that imposes the encoding phase pattern $\theta(\omega)$. The encoded photon propagates through an optical fiber channel to a receiving station implementing a decoding mask $D(\omega)=e^{i\tilde{\theta}(\omega)}$. The decoded photon then interacts with a two-level emitter such as a trapped atom, ion, or quantum dot coupled to a single-mode waveguide or high-NA free-space interface. The excitation probability is measured via fluorescence detection or by monitoring changes in transmission.
Systematic variations of the code length $N_{0}$, phase-error amplitude $\sigma_{\phi}$, and decoding alignment allow direct experimental verification of (i) the exponential sensitivity $P_{e}\propto e^{-\sigma^{2}_{\phi}}$ predicted in Section IV B, (ii) the $N_{0}^{-1}$ scaling of cross-talk between multiple users (Section IV C), and (iii) the overall dependence of $P_{e}$ on $W/\gamma$ and $\beta$. Because all components-pulse shapers, fiber channels, and single emitters-are compatible with current laboratory technology, this platform-agnostic implementation offers a clear route for testing the theoretical predictions and for deploying spectral-encoded quantum communication protocols based on the principles developed in this study.
V Conclusion
We analyzed the interaction between a two-level atom and a spectrally encoded single-photon wave packet using the Heisenberg-Langevin approach. The results show that spectral phase encoding, while unitary, reshapes the photon’s temporal profile and reduces its peak intensity at the atom, thereby lowering the excitation probability compared with an unencoded pulse. The efficiency is governed by the temporal-mode overlap between the incident photon and the atom’s optimal absorption mode, incomplete decoding or phase errors directly reduce this overlap.
From a network perspective, the atom behaves as a mode-selective and phase-sensitive receiver. Perfect conjugate decoding restores the original mode and strong coupling, whereas mismatched codes are naturally rejected, providing a built-in spectral filter for secure or multiplexed communication.
Efficient operation requires careful control of spectral bandwidth, phase stability, mode matching, and detuning compensation. These parameters jointly determine the achievable overlap and thus the coupling efficiency. The framework presented here offers a compact description linking atomic excitation to a spectral-temporal overlap functional, supplying both a diagnostic for performance degradation under encoding and practical guidelines for implementing spectrally encoded photon-atom interfaces in scalable quantum networks.